OpenStudy (anonymous):

add and simplify t+2/3t+t-4/6

OpenStudy (anonymous):

(6/6) t+2/3t + t-4/6 (3t/3t)

OpenStudy (anonymous):

\[\frac{ 6t+12+3t^2-12t }{ 3t(6) }\]

OpenStudy (anonymous):

\[\frac{ 3t^2-6t+12 }{ 3(t+2) }\]

OpenStudy (anonymous):

\[\frac{ 3(t^2-2t+4) }{ 3(t-2) }\]

OpenStudy (anonymous):

\[\frac{ t^2-2t+4 }{ t-2 }\]

OpenStudy (anonymous):

since you can't factor 3(t^2-2t+4)